When a small sphere moves at a low speed through a fluid, the viscous force F opposing the motion is found experimentally to depend on the radius r, velocity v, and the coefficient of viscosity n of the fluid. Find the force dimension of the force F
1
Expert's answer
2019-12-19T11:37:20-0500
"F=r^av^b\eta^c"

"\dim [F]=(\dim [r])^a(\dim [v])^b(\dim [\eta])^c"

"MLT^{-2}=(L)^a(LT^{-1})^b(ML^{-1}T^{-1})^c"

"c=1"

"-2=-b-1\to b=1"

"1=a+1-1\to a=1"

Thus,


"F=\eta rv"